# Understanding the Limits of Deep Tabular Methods with Temporal Shift

### v1

@@ Abstract
Deep tabular models have demonstrated remarkable success on i.i.d. data, excelling in a variety of structured data tasks. However, their performance often deteriorates under temporal distribution shifts, where trends and periodic patterns are present in the evolving data distribution over time. In this paper, we explore the underlying reasons for this failure in capturing temporal dependencies. We begin by investigating the training protocol, revealing a key issue in how model selection perform. While existing approaches use temporal ordering for splitting validation set, we show that even a random split can significantly improve model performance. By minimizing the time lag between training data and test time, while reducing the bias in validation, our proposed training protocol significantly improves generalization across various methods. Furthermore, we analyze how temporal data affects deep tabular representations, uncovering that these models often fail to capture crucial periodic and trend information. To address this gap, we introduce a plug-and-play temporal embedding method based on Fourier series expansion to learn and incorporate temporal patterns, offering an adaptive approach to handle temporal shifts. Our experiments demonstrate that this temporal embedding, combined with the improved training protocol, provides a more effective and robust framework for learning from temporal tabular data.
@@ 1 Introduction
Tabular data is one of the most prevalent data formats in a wide range of real-world applications, such as e-commerce (Nederstigt et al., 2014) and healthcare (Hassan et al., 2020). It consists of instances (rows) and features (columns), where the label can either be categorical or continuous, corresponding to classification and regression tasks, respectively. The goal is to learn a mapping strategy from features to labels that can be directly applied to independent and identically distributed (i.i.d.) test data for accurate predictions (Bishop, 2007; Mohri et al., 2012). While tree-based methods remain powerful (Breiman, 2001; Chen & Guestrin, 2016; Ke et al., 2017; Prokhorenkova et al., 2018), recent advancements in deep learning methods for tabular data have demonstrated promising results (Klambauer et al., 2017; Wang et al., 2021b; Gorishniy et al., 2021, 2024, 2025; Hollmann et al., 2023, 2025; Holzmüller et al., 2024; Ye et al., 2023, 2025; Liu & Ye, 2025).
Most machine learning approaches are built on the assumption of i.i.d. data. However, in open environments (Zhou, 2022), distribution shifts (Gardner et al., 2023; Tschalzev et al., 2024) frequently occur between the training and testing data, which can manifest as shifts in the feature space, label space, or the mapping between them. Moreover, in practical applications, data often exhibits temporal distribution shifts (Rubachev et al., 2025), a particularly common type of distribution shift that introduces additional challenges: instead of just shifts between training and test sets, temporal shifts can occur within the training set or the test set itself. For example, in house price prediction (Matveev & Sidorova, 2017), the task is to predict future house prices using historical transaction data, which includes features such as location, neighborhood conditions, and economic factors. However, in actual practice, the trend of housing policies or the periodic fluctuations of public sentiment can also play crucial roles in influencing house prices.
@@
Since temporal shifts are common and present significant challenges, we have turned our attention on whether tabular data models that perform well in classic i.i.d. scenarios can effectively manage temporal shifts. In recent studies, Rubachev et al. (2025) introduced the TabReD benchmark. By comparing the performance differences of various methods on the original temporal shift dataset and the i.i.d. dataset constructed through shuffling, their observations revealed that while models with MLP architectures exhibit relative robustness during temporal shifts, retrieval-based methods, which are highly competitive in current benchmarks (Ye et al., 2024), suffer a marked performance degradation in temporal shift scenarios. This observation demonstrates that the occurrence of temporal shifts can introduce significant biases in the evaluation of model performance, further emphasizing the importance of understanding and managing these temporal shifts.
@@
In temporal shift tasks, the absence of accurate test data validation during the training stage (Blanchard et al., 2021) makes model selection more impactful, as deep learning models are optimized epoch-wise. TabReD employs a temporal splitting strategy to match the test scenario, utilizing earlier data for training and later data prior to as the validation set for model selection. Surprisingly, we discovered that by merely altering the splitting strategy on the same training data, even when randomly splitting the training and validation sets and ignoring temporal order, the model outperformed the temporal split, resulting in significantly enhanced performance, as illustrated in Figure 2. Moreover, the performance rankings of the models became more consistent with established results on i.i.d. data, with retrieval-based methods showing more distinct improvements. This unexpected finding underscores the urgent need to investigate effective data splitting strategies to accurately assess model performance and reveal its true capabilities in the presence of temporal shifts. Despite the performance improvement, the random split exhibits pronounced instability in temporal shift context, which also requires attention.
@@
We begin by analyzing the factors contributing to the ineffectiveness of temporal splitting in this context. While temporal-based splitting is commonly employed in forecasting tasks to maintain causal relationships (Bergmeir & Benítez, 2012), the most intuitive difference when adopting a temporal split in tabular learning lies in the time lag between the training and test sets, and the bias in validation, since the data closest to the test time are not used for training, and the shift degree of test set relative to validation set is more significant in temporal splits. By investigating the impact of these two factors, we find that minimizing the training lag concentrates the model’s performance closer to the test time, while reducing validation bias by aligning the shift degree makes the model better generalize on test data. Building on these insights, we propose a training protocol along with an improved temporal split that leverages these advantages, resulting in a comparable performance to random splitting while providing better stability.
@@ 2.1 Tabular Machine Learning
Tabular data is a common format across various applications, and the main solutions can be categorized into tree-based methods, token-based methods, retrieval-based methods, ensemble-based methods, and MLP architecture methods. Classical tree-based methods like Random Forest (Breiman, 2001), XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018), remain powerful and widely adopted in real-world scenarios. FT-Transformer (FT-T) (Gorishniy et al., 2021) is a token-based method that utilizes the transformer architecture, TabR (Gorishniy et al., 2024) and ModernNCA (Ye et al., 2025) are retrieval-based methods that make predictions by retrieving neighbors in the representation space, TabM (Gorishniy et al., 2025) provides an ensemble-based method on MLP, while other methods like SNN (Klambauer et al., 2017), DCNv2 (Wang et al., 2021b) and MLP-PLR (Gorishniy et al., 2022) focus on improving the MLP architecture. With the continuous improvement of deep tabular models on established benchmarks (McElfresh et al., 2023; Ye et al., 2024), the practical deployment of such models has become a pressing consideration.
@@ 2.2 Distribution Shift in Tabular Data
Current research on distribution shifts can be broadly categorized into two main approaches. The first category focuses on scenarios in which target domain data is partially available. In these cases, transfer learning techniques are commonly employed to dynamically adjust model parameters during deployment by leveraging test-time data. TableShift (Gardner et al., 2023) applies various classical methods of domain generalization (Ganin et al., 2016) and domain adaptation (Sun & Saenko, 2016) into deep tabular learning, alongside the recently proposed test-time adaptation techniques for tabular data (Kim et al., 2024). While effective in certain contexts, these approaches often assume the availability of target domain information at test time, which may be infeasible in real-world settings. In our experimental setup, the test time information is entirely invisible during both the training and testing stage, thus methods in this category are not applicable to our setting.
The second category addresses situations in which target domain data is entirely unavailable, representing a more common and challenging scenario. Approaches within this category can be further divided into two types: those aimed at enhancing model robustness and those focused on the active learning of shift patterns. The first type seeks to improve the inherent robustness and generalization of models, thereby indirectly mitigating the impact of distribution shifts. For instance, Gorishniy et al. (2025) demonstrates the effectiveness of ensemble strategies in addressing distribution shifts in tabular data. Our exploration of the training protocol serves as an effective approach to enhancing model generalization performance. The second type incorporates knowledge of distribution shifts directly into the model through adaptive methods. One such approach is presented by Helli et al. (2024), which employs second-order models to capture and adapt to shifts based on learned causal relationships. However, methods in this category are heavily dependent on specific model architectures, such as PFN (Hollmann et al., 2023), and tend to be computationally expensive. Consequently, we did not include a comparison with this category of methods. In contrast, our temporal embedding method offers a lightweight and plug-and-play solution for achieving temporal adaptability.
While TableShift (Gardner et al., 2023) emphasizes domain-to-domain shifts, TabReD (Rubachev et al., 2025) introduces the concept of temporal shifts. They argue that all tabular data is inherently temporal, advocating for the use of temporal splits, especially in industrial applications where data is typically feature-rich and includes visible timestamps. In this study, we further investigate a training protocol and propose a more effective and robust framework for learning from temporal tabular data.
@@ 2.3 Distribution Shift in Other Domains
The study of distribution shifts originated in computer vision, with early research primarily focusing on transfer learning techniques to address domain-to-domain shifts, including domain generalization (Blanchard et al., 2011, 2021), domain adaptation (Ganin & Lempitsky, 2015), and test-time adaptation (Wang et al., 2021a). Later, the focus expanded to encompass continual distribution shifts, with strategies for adapting models to sequential domain changes (Wang et al., 2022), as well as addressing recurring (Hoang et al., 2024) and non-i.i.d. (Gong et al., 2022) temporal shifts. Wild-Time (Yao et al., 2022) explores real-world temporal shifts but translates them into domain-to-domain settings, neglecting temporal continuity. In comparison, these methods are primarily designed for images and most utilize transfer learning. Many image-based methods, when directly applied to tabular data, are considered to perform poorly (Gardner et al., 2023). This may be due to the more complex and diverse distribution shifts in tabular data, which involve greater temporal dependencies and continuity.
@@ 3 Preliminary
The objective function can be decomposed into two components: , where for feature extraction and for prediction. In MLP architectures, both and are neural networks, with consisting of multiple layers and being the final output layer. In ensemble-based methods, multiple models are trained, and the final prediction is obtained by averaging the outputs, which reduces variance and improves generalization. Retrieval-based methods use a neural network for and a non-parametric prediction method, such as a soft KNN adopted in ModernNCA (Ye et al., 2025), for , where predictions are based on the closest neighbors of “candidates” — instances from the training set that are mapped to the representation space and serve as potential reference points during prediction.
A temporal tabular dataset collected before for model training and deployment can be represented as , where denotes the set of instances collected at time , each attached with its timestamp. is the time at which training is performed. After training, the model is deployed to an open environment and evaluated on . The training data need to be further split into training and validation set in training stage. TabReD (Rubachev et al., 2025) adopts a temporal split where the data is divided at , such that and .
A temporal distribution shift refers to a specific type of distribution shift where the data is collected sequentially over time, and the underlying data distribution evolves as time progresses. Formally, let , denote the feature space and label space at time , respectively. A temporal distribution shift concurrently suffer from covariant shift, label shift, and concept shift, i.e., , , or for some .
@@
4 Why Temporal Split is Ineffective for Tabular Data?
The fundamental challenge in addressing temporal shifts stems from the inherent difficulty in characterizing evolving data distributions, which requires models to maintain robustness against unknown future variations. Temporal splits are frequently used in forecasting-related tasks to uphold the sequential dependencies inherent in the data (Bergmeir & Benítez, 2012; Zeng et al., 2023; Han et al., 2024), yet they generally show inferior performance compared to random splits in tabular data contexts, as evidenced in Section 1 and Figure 2. This counterintuitive phenomenon warrants systematic analysis through multiple perspectives: We first start with the most intuitive distinction between random split and temporal split, and verify the hypothesis of how training lag (Section 4.1) and validation bias (Section 4.2) impact the model performance. Then we discover the equivalence of the validation set in different temporal directions (Section 4.3). Building on this, we propose an enhanced splitting strategy (Section 4.4) for temporally shifted tabular data, which achieves performance comparable to random splitting while offering improved stability.
@@ 4.1 Training Lag
From the perspective of the training set, temporal split adopted in Rubachev et al. (2025) introduces a temporal lag between training and test set, while random split provides more instances closer to the test time for training. An intuitive hypothesis is that instances closer to the test time are more reliable, as the distribution near the test time tends to be more similar to the actual test-time distribution, and using them for training could have a greater effect than for validation. Hence we adapt a pair of splitting strategies, shown in Figure 3 left (a) and (b), where identical validation and test sets are maintained while varying the lag between training and test set, thereby isolating and quantifying the impact of training lag on model performance.
@@ 4.2 Validation Bias
Current deep tabular methods rely on information from the validation set for model selection, as deep learning models are optimized epoch-wise during training. In the context of temporal shifts, this reliance becomes particularly problematic, since there is a lack of accurate validation of the test data in the training stage (Ganin & Lempitsky, 2015; Blanchard et al., 2021). The time gap between the training and test sets is larger than the gap between the training and validation sets in the previous temporal split, leading to a more significant distribution shift at test time. This makes it more challenging to accurately predict test-time instances compared to validation, thereby introducing bias into the validation process. Therefore we further design the splitting strategy shown in Figure 3 left (c), which shares the training and test sets with (a), but the latter has a more considerable validation bias since the time interval difference between train-val and train-test is much larger.
@@ 4.3 Bridging the Past and Future in Temporal Split
To illustrate how reducing the training lag and validation bias improves model performance, we visualize the loss distribution of the model across different time slices under different validation splitting strategies. Figure 4 left shows the loss distribution for the Cooking Time dataset using MLP-PLR model, it clearly shows how reducing training lag and validation bias improves the model performance. Each splits achieve the best performance among the training-available time slices, which allows the model to perform better on instances closer to the test time in the non-lagged split (b) compared to the lagged splits (a) and (c). Furthermore, the higher-biased split (c) works better on training splits but struggles on the test splits, while the lower-biased split (a) achieves a more balanced performance across training and test splits. This indicates that a lower-biased validation set brings a more precise direction of the test-time distribution, enabling the model to generalize better. The above observation again enhances the insights of training lag and validation bias. On a deeper level, reducing the training lag concentrates the model’s performance around a time point closer to the test time, while precise validation ensures that the model’s performance can effectively generalize from this concentrated time point to the test period.
In addition, another interesting observation in Figure 4 left is that the model in split (b) never meet the most former data splits during training and validation, but achieves a relatively well performance on it compared to the test splits. Building on the insight that the validation set is mainly for maintaining generalization performance, this finding inspired us to explore whether data from the opposite temporal direction could be an effective validation set, which is rely on the assumption that the temporal shifts distributed uniformly across each time slices. We further use a heatmap to visualize the distribution distance between different time slices using Maximum Mean Discrepancy (MMD) (Gretton et al., 2006), as shown in Figure 4 right. The MMD heatmap reveals that all datasets exhibit temporal shifts characterized by trends and multiple periodic components. The regular diagonal stripes both indicate the presence of periodicity, and suggest that the sample distributions at identical time intervals are similar, which confirms the empirical uniformity of temporal shifts across time slices.

### v2

@@ Abstract
Deep tabular models have demonstrated remarkable success on i.i.d. data, excelling in a variety of structured data tasks. However, their performance often deteriorates under temporal distribution shifts, where trends and periodic patterns are present in the evolving data distribution over time. In this paper, we explore the underlying reasons for this failure in capturing temporal dependencies. We begin by investigating the training protocol, revealing a key issue in how model selection performs. While existing approaches use temporal ordering for splitting validation set, we show that even a random split can significantly improve model performance. By minimizing the time lag between training data and test time, while reducing the bias in validation, our proposed training protocol significantly improves generalization across various methods. Furthermore, we analyze how temporal data affects deep tabular representations, uncovering that these models often fail to capture crucial periodic and trend information. To address this gap, we introduce a plug-and-play temporal embedding method based on Fourier series expansion to learn and incorporate temporal patterns, offering an adaptive approach to handle temporal shifts. Our experiments demonstrate that this temporal embedding, combined with the improved training protocol, provides a more effective and robust framework for learning from temporal tabular data.
@@ 1 Introduction
Tabular data is one of the most prevalent data formats in a wide range of real-world applications, such as e-commerce (nederstigt2014floppies) and healthcare (hassan2020machine). It consists of instances (rows) and features (columns), where the label can either be categorical or continuous, corresponding to classification and regression tasks, respectively. The goal is to learn a mapping strategy from features to labels that can be directly applied to independent and identically distributed (i.i.d.) test data for accurate predictions (bishop2007pattern; mohri2012foundations). While tree-based methods remain powerful (breiman2001random; chen2016xgboost; ke2017lightgbm; prokhorenkova2018catboost), recent advancements in deep learning methods have demonstrated promising results (klambauer2017self; arik2021tabnet; wang2021dcn; gorishniy2021revisiting; gorishniy2024tabr; gorishniy2024tabm; hollmann2022tabpfn; hollmann2025accurate; ye2023rethinkingpretrainingtabulardata; ye2024modern; holzmuller2024better; jiang2024tabular; jiang2025representation; liu2025tabpfnunleashed; liu2025makestillfurtherprogress).
Most machine learning approaches are built on the assumption of i.i.d. data. However, in open environments (zhou2022open), distribution shifts (gardner2023benchmarking; tschalzev2024data) frequently occur between the training and testing data, which can manifest as shifts in the feature space, label space, or the mapping between them. Moreover, in practical applications, data often exhibits temporal distribution shifts (rubachev2024tabred), a particularly common type of distribution shift that introduces additional challenges: instead of just shifts between training and test sets, temporal shifts can occur within the training set or the test set itself. For example, in house price prediction (matveev2017sberbank), the task is to predict future house prices using historical transaction data, which includes features such as location, neighborhood conditions, and economic factors. However, in actual practice, the trend of housing policies or the periodic fluctuations of public sentiment can also play crucial roles in influencing house prices.
@@
Since temporal shifts are common and present significant challenges, we have turned our attention on whether tabular data models that perform well in classic i.i.d. scenarios can effectively manage temporal shifts. In recent studies, rubachev2024tabred introduced the TabReD benchmark. By comparing the performance differences of various methods on the original temporal shift dataset and the i.i.d. dataset constructed through shuffling, their observations revealed that while models with MLP architectures exhibit relative robustness during temporal shifts, retrieval-based methods, which are highly competitive in current benchmarks (ye2024closer), suffer a marked performance degradation in temporal shift scenarios. This observation demonstrates that the occurrence of temporal shifts can introduce significant biases in the evaluation of model performance, further emphasizing the importance of understanding and managing these temporal shifts.
@@
In temporal shift tasks, the absence of accurate test data validation during the training stage (blanchard2021domain) makes model selection more impactful, as deep learning models are optimized epoch-wise. TabReD employs a temporal splitting strategy to match the test scenario, utilizing earlier data for training and later data prior to as the validation set for model selection. Surprisingly, we discovered that by merely altering the splitting strategy on the same training data, even when randomly splitting the training and validation sets and ignoring temporal order, the model outperformed the temporal split, resulting in significantly enhanced performance, as illustrated in Figure 2. Moreover, the performance rankings of the models became more consistent with established results on i.i.d. data, with retrieval-based methods showing more distinct improvements. This unexpected finding underscores the urgent need to investigate effective data splitting strategies to accurately assess model performance and reveal its true capabilities in the presence of temporal shifts. Despite the performance improvement, the random split exhibits pronounced instability in temporal shift context, which also requires attention.
@@
We begin by analyzing the factors contributing to the ineffectiveness of temporal splitting in this context. While temporal-based splitting is commonly employed in forecasting tasks to maintain causal relationships (bergmeir2012use), the most intuitive difference when adopting a temporal split in tabular learning lies in the time lag between the training and test sets, and the bias in validation, since the data closest to the test time are not used for training, and the shift degree of test set relative to validation set is more significant in temporal splits. By investigating the impact of these two factors, we find that minimizing the training lag concentrates the model’s performance closer to the test time, while reducing validation bias by aligning the shift degree makes the model better generalize on test data. Building on these insights, we propose a training protocol along with an improved temporal split that leverages these advantages, resulting in a comparable performance to random splitting while providing better stability.
@@ 2.1 Tabular Machine Learning
Tabular data is a common format across various applications, and the main solutions can be categorized into tree-based methods, token-based methods, retrieval-based methods, ensemble-based methods, and MLP architecture methods. Classical tree-based methods like Random Forest (breiman2001random), XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), and CatBoost (prokhorenkova2018catboost), remain powerful and widely adopted in real-world scenarios. FT-Transformer (FT-T) (gorishniy2021revisiting) is a token-based method that utilizes the transformer architecture, TabR (gorishniy2024tabr) and ModernNCA (ye2024modern) are retrieval-based methods that make predictions by retrieving neighbors in the representation space, TabM (gorishniy2024tabm) provides an ensemble-based method on MLP, while other methods like SNN (klambauer2017self), DCNv2 (wang2021dcn) and MLP-PLR (gorishniy2022embeddings) focus on improving the MLP architecture. With the continuous improvement of deep tabular models on established benchmarks (mcelfresh2023when; ye2024closer), the practical deployment of such models has become a pressing consideration.
@@ 2.2 Distribution Shift in Tabular Data
Current research on distribution shifts can be broadly categorized into two main approaches. The first category focuses on scenarios in which target domain data is partially available. In these cases, transfer learning techniques are commonly employed to dynamically adjust model parameters during deployment by leveraging test-time data. TableShift (gardner2023benchmarking) applies various classical methods of domain generalization (ganin2016domain) and domain adaptation (sun2016deepcoral) into deep tabular learning, alongside the recently proposed test-time adaptation techniques for tabular data (kim2024adaptable). While effective in certain contexts, these approaches often assume the availability of target domain information at test time, which may be infeasible in real-world settings. In our experimental setup, the test time information is entirely invisible during both the training and testing stage, thus methods in this category are not applicable to our setting.
The second category addresses situations in which target domain data is entirely unavailable, representing a more common and challenging scenario. Approaches within this category can be further divided into two types: those aimed at enhancing model robustness and those focused on the active learning of shift patterns. The first type seeks to improve the inherent robustness and generalization of models, thereby indirectly mitigating the impact of distribution shifts. For instance, gorishniy2024tabm demonstrates the effectiveness of ensemble strategies in addressing distribution shifts in tabular data. Our exploration of the training protocol serves as an effective approach to enhancing model generalization performance. The second type incorporates knowledge of distribution shifts directly into the model through adaptive methods. One such approach is presented by helli2024drift, which employs second-order models to capture and adapt to shifts based on learned causal relationships. However, methods in this category are heavily dependent on specific model architectures, such as PFN (hollmann2022tabpfn), and tend to be computationally expensive. Consequently, we did not include a comparison with this category of methods. In contrast, our temporal embedding method offers a lightweight and plug-and-play solution for achieving temporal adaptability.
While TableShift (gardner2023benchmarking) emphasizes domain-to-domain shifts, TabReD (rubachev2024tabred) introduces the concept of temporal shifts. They argue that all tabular data is inherently temporal, advocating for the use of temporal splits, especially in industrial applications where data is typically feature-rich and includes visible timestamps. In this study, we further investigate a training protocol and propose a more effective and robust framework for learning from temporal tabular data.
@@ 2.3 Distribution Shift in Other Domains
The study of distribution shifts originated in computer vision, with early research primarily focusing on transfer learning techniques to address domain-to-domain shifts, including domain generalization (blanchard2011generalizing; blanchard2021domain), domain adaptation (ganin2015unsupervised), and test-time adaptation (wang2021tent). Later, the focus expanded to encompass continual distribution shifts, with strategies for adapting models to sequential domain changes (wang2022continual), as well as addressing recurring (hoang2024persistent) and non-i.i.d. (gong2022note) temporal shifts. Wild-Time (yao2022wild) explores real-world temporal shifts but translates them into domain-to-domain settings, neglecting temporal continuity. In comparison, these methods are primarily designed for images and most utilize transfer learning. Many image-based methods, when directly applied to tabular data, are considered to perform poorly (gardner2023benchmarking). This may be due to the more complex and diverse distribution shifts in tabular data, which involve greater temporal dependencies and continuity.
@@ 3 Preliminary
The objective function can be decomposed into two components: , where for feature extraction and for prediction. In MLP architectures, both and are neural networks, with consisting of multiple layers and being the final output layer. In ensemble-based methods, multiple models are trained, and the final prediction is obtained by averaging the outputs, which reduces variance and improves generalization. Retrieval-based methods use a neural network for and a non-parametric prediction method, such as a soft -NN adopted in ModernNCA (ye2024modern), for , where predictions are based on the closest neighbors of “candidates” — instances from the training set that are mapped to the representation space and serve as potential reference points during prediction.
A temporal tabular dataset collected before for model training and deployment can be represented as , where denotes the set of instances collected at time , each attached with its timestamp. is the time at which training is performed. After training, the model is deployed to an open environment and evaluated on . The training data need to be further split into training and validation set in training stage. TabReD (rubachev2024tabred) adopts a temporal split where the data is divided at , such that and .
A temporal distribution shift refers to a specific type of distribution shift where the data is collected sequentially over time, and the underlying data distribution evolves as time progresses. Formally, let and denote the feature space and label space at time , respectively. For and , a temporal distribution shift may involve changes in , , , or even or , for some .
@@
4 Why Temporal Split Ineffective?
The fundamental challenge in addressing temporal shifts stems from the inherent difficulty in characterizing evolving data distributions, which requires models to maintain robustness against unknown future variations. Temporal splits are frequently used in forecasting-related tasks to uphold the sequential dependencies inherent in the data (bergmeir2012use; zeng2023transformers; han2024revisiting), yet they generally show inferior performance compared to random splits in tabular data contexts, as evidenced in Section 1 and Figure 2. This counterintuitive phenomenon warrants systematic analysis through multiple perspectives: We first start with the most intuitive distinction between random split and temporal split, and verify the hypothesis of how training lag (Section 4.1) and validation bias (Section 4.2) impact the model performance. Then we discover the equivalence of the validation set in different temporal directions (Section 4.3). Building on this, we propose an enhanced splitting strategy (Section 4.4) for temporally shifted tabular data, which achieves performance comparable to random splitting while offering improved stability.
@@ 4.1 Training Lag
From the perspective of the training set, temporal split adopted in rubachev2024tabred introduces a temporal lag between training and test set, while random split provides more instances closer to the test time for training. An intuitive hypothesis is that instances closer to the test time are more reliable, as the distribution near the test time tends to be more similar to the actual test-time distribution, and using them for training could have a greater effect than for validation. Hence we adapt a pair of splitting strategies, shown in Figure 3 left (a) and (b), where identical validation and test sets are maintained while varying the lag between training and test set, thereby isolating and quantifying the impact of training lag on model performance.
@@ 4.2 Validation Bias
Current deep tabular methods rely on information from the validation set for model selection, as deep learning models are optimized epoch-wise during training. In the context of temporal shifts, this reliance becomes particularly problematic, since there is a lack of accurate validation of the test data in the training stage (ganin2015unsupervised; blanchard2021domain). The time gap between the training and test sets is larger than the gap between the training and validation sets in the previous temporal split, leading to a more significant distribution shift at test time. This makes it more challenging to accurately predict test-time instances compared to validation, thereby introducing bias into the validation process. Therefore we further design the splitting strategy shown in Figure 3 left (c), which shares the training and test sets with (a), but the latter has a more considerable validation bias since the time interval difference between train-val and train-test is much larger.
@@ 4.3 Bridging the Past and Future in Temporal Split
To illustrate how reducing the training lag and validation bias improves model performance, we visualize the loss distribution of the model across different time slices under different validation splitting strategies. Figure 4 left shows the loss distribution for the CT dataset using MLP-PLR model, it clearly shows how reducing training lag and validation bias improves the model performance. Each splits achieve the best performance among the training-available time slices, which allows the model to perform better on instances closer to the test time in the non-lagged split (b) compared to the lagged splits (a) and (c). Furthermore, the higher-biased split (c) works better on training splits but struggles on the test splits, while the lower-biased split (a) achieves a more balanced performance across training and test splits. This indicates that a lower-biased validation set brings a more precise direction of the test-time distribution, enabling the model to generalize better. The above observation again enhances the insights of training lag and validation bias. On a deeper level, reducing the training lag concentrates the model’s performance around a time point closer to the test time, while precise validation ensures that the model’s performance can effectively generalize from this concentrated time point to the test period.
In addition, another interesting observation in Figure 4 left is that the model in split (b) never meet the most former data splits during training and validation, but achieves a relatively well performance on it compared to the test splits. Building on the insight that the validation set is mainly for maintaining generalization performance, this finding inspired us to explore whether data from the opposite temporal direction could be an effective validation set, which is rely on the assumption that the temporal shifts distributed uniformly across each time slices. We further use a heatmap to visualize the distribution distance between different time slices using Maximum Mean Discrepancy (MMD) (gretton2006kernel), as shown in Figure 4 right. The MMD heatmap reveals that all datasets exhibit temporal shifts characterized by trends and multiple periodic components. The regular diagonal stripes both indicate the presence of periodicity, and suggest that the sample distributions at identical time intervals are similar, which confirms the empirical uniformity of temporal shifts across time slices.